OpenStudy (anonymous):

the sum of three consecutive number is 84 . what is the largest of the three number?

OpenStudy (anonymous):

Consecutive numbers can be written as: x, x+1, x+2,... and so on for as many as you need. So for this problem, we only need the first three. And their SUM is 84. So we can write this equation: x + x+1 + x+2 = 84 Be careful though: solving for x IS NOT the answer to this problem. After you solve for x, you have to plug back in to find the largest of the 3 numbers, which is represented by x+2. So solve that equation for x, then plug into x + 2 to find the largest of the three numbers.

OpenStudy (anonymous):

dont get it can you show work n answer to help me

OpenStudy (anonymous):

First number: x Second number: x+1 Third number: x+2 So to add those three numbers, you get: x + x + 1 + x + 2. They tell you the sum is 84, so you build this equation: \[x+x+1+x+2=84\] Combine like terms next. There are 3 single x's, so that's 3x. 1+2 is 3. So: \[3x+3=84\] Can you solve that equation for x? It's a two-step equation.

OpenStudy (anonymous):